Troy Engines, Ltd., manufactures a variety of engines for use in heavy equipment. The company has always produced all of the necessary parts for its engines, including all of the carburetors. An outside supplier has offered to sell one type of carburetor to Troy Engines, Ltd., for a cost of $33 per unit. To evaluate this offer, Troy Engines, Ltd., has gathered the following information relating to its own cost of producing the carburetor internally: 18,000 Units Per Unit Year Per Direct materials Direct labor Variable manufacturing overhead Fixed manufacturing overhead, traceable Fixed manufacturing overhead, allocated Total cost 15 270,000 9 162,000 72,000 6 108,000 9 162, 000 43 $ 774,000 One-third supervisory salaries; two-thirds depreciation of special equipment (no resale value) Required 1. Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 18,000 carburetors from the outside supplier? 2. Should the outside suppliers offer be accepted? 3. Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $180,000 per year. Given this new assumption, what would be financial advantage (disadvantage) of buying 18,000 carburetors from the outside supplier? 4. Given the new assumption in requirement 3, should the outside suppliers offer be accepted?

Complete this question by entering your answers in the tabs below. Required 1Required 2Required 3 Required 4 Assuming the company has no alternative use for the facilities that are now being used to produce the carburetors, what would be the financial advantage (disadvantage) of buying 18,000 carburetors from the outside supplier?

Complete this question by entering your answers in the tabs below. Required 1 Required 2 Required 3Required 4 Suppose that if the carburetors were purchased, Troy Engines, Ltd., could use the freed capacity to launch a new product. The segment margin of the new product would be $180,000 per year. Given this new assumption, what would be the financial advantage (disadvantage) of buying 18,000 carburetors from the outside supplier? K Required 2 Required4>
Answer #1

1) Differential analysis

Make Buy
Direct material 270000
Direct labor 162000
Variable manufacturing overhead 72000
Fixed manufacturing overhead (108000/3) 36000
Purchase cost (18000*33) 594000
Total relevant cost 540000 594000

Financial (disadvantage) = 540000-594000 = -$54000

2) No should not accept the offer

3) Differential analysis

Make Buy
Direct material 270000
Direct labor 162000
Variable manufacturing overhead 72000
Fixed manufacturing overhead (108000/3) 36000
Opportunity Cost 180000
Purchase cost (18000*33) 594000
Total relevant cost 720000 594000

Financial advantage = 720000-594000 = -$126000

4)yes should accept the offer